OpenStudy (anonymous):

5(x-8)+8=5x-8(8-x) x^2-x=?

OpenStudy (***[isuru]***):

hi, 5(x-8)+8=5x-8(8-x) 5(x-8)+8=5x+8(x-8) 5(x-8) -8(x-8) =5x -8 (x -8)(5-8) = 5x - 8 8-5x =3(x-8) 8x -16 = 0 Now u can follow these steps Or r u looking for a way to deduce the answer?